### v2

@@ 5 Beyond basic variational-like IMs
Various tweaks to the basic procedure described in Section 3 might be considered in order applications to speed up computations. If there is a computational bottleneck remaining in the proposal from Section 3, then it would have to be that the IM’s possibility contour must be evaluated at $M$-many points in each iteration of the stochastic approximation. While this is not impractically expensive in some applications, including those presented above, it could be more of a burden in other applications. A related challenge is that a scalar index $\xi$ on the variational family, as we have focused on exclusively so far, may not be sufficiently flexible. Increasing the flexibility by considering higher-dimensional $\xi$ would likewise increase the computational burden, so care is needed. Here we aim to address both of the aforementioned challenges.
@@
That is, the components of $\xi$ act as multipliers on the singular values of $J_{n}^{-1}$. This indeed generalizes our previously-considered Gaussian approximation family since, if $\xi$ is a scalar, then we recover the previous covariance matrix: $J_{n}(\xi)=J_{n}/\xi^{2}$. Let $K_{n,\alpha}^{\xi}$ denote the $100(1-\alpha)$credible set corresponding to the distribution $\mathsf{Q}_{n,\alpha}^{\xi}\in\mathscr{Q}^{\text{var}}$. In the present $d$-dimensional Gaussian case, this is an ellipsoid in $\mathbb{R}^{d}$, i.e.,
$K_{n,\alpha}^{\xi}=\{\theta\in\mathbb{R}^{d}:(\theta-\hat{\theta}_{n})^{\top}J_{n}(\xi)\,(\theta-\hat{\theta}_{n})\leq\chi_{d}^{2}(1-\alpha)\},$
@@
$K_{n,\alpha}^{\xi}\supseteq C_{\alpha}(x^{n}),$
@@
$\sup_{\theta\not\in K_{n,\alpha}^{\xi}}\pi_{n}(\theta)\leq\alpha.$
For the Gaussian family of approximations we are considering here, the mode of $\pi_{n}$, the maximum likelihood estimator, is in the interior of each credible set, so the behavior of $\pi_{n}$ outside $K_{n,\alpha}^{\xi}$ is determined by its behavior on the boundary. Moreover, since equality in the above display implies a near-perfect match between the IM’s confidence sets and the posited Gaussian credible sets, a reasonable goal is to find a root to the function
$g_{\alpha}(\xi):=\max_{\theta\in\text{bdry}(K_{n,\alpha}^{\xi})}\pi_{n}(\theta)-\alpha,$
where $\text{bdry}(A)$ denotes the boundary of a set $A$. In designing an iterative algorithm that makes steps towards this root, it is desirable to allow for steps of different magnitudes depending on the direction, and this requires care. For our present Gaussian approximation family, a natural—albeit minimal—representation of the boundary of the credible region $K_{n,\alpha}^{\xi}$ is given by the collection of vectors
@@
At least intuitively, $\hat{g}_{\alpha,s}(\xi)$ being less than or greater than 0 determines if the credible set $K_{n,\alpha}^{\xi}$ is too large or too small, respectively in the $u_{s}$-direction. From here, we can apply the same stochastic approximation procedure to determine a sequence $(\xi^{(t)}:t\geq 1)$, this time of $d$-vectors, i.e.,
@@
that converges to the root of $\hat{g}_{\alpha}$ which we expect to be close to the root of $g_{\alpha}$. The details of this new but familiar strategy are presented in Algorithm 2. The chief advantage of this new strategy compared to that summarized in Algorithm 1 is computational savings: each iteration of stochastic approximation in Algorithm 2 only requires (Monte Carlo) evaluation of the IM contour at $2d$ many points, which would typically be far fewer than then number of Monte Carlo samples $M$ in Algorithm 1 needed to get a good approximation of $\mathsf{Q}_{n,\alpha}^{\xi}\{\pi_{n}(\Theta)>\alpha\}$.
@@
We will illustrate this new version of the Gaussian variational family and approximation algorithm with two examples. The first is in the relatively simple two-parameter gamma model; the second is a relatively high-dimensional model involving a penalty, intended to serve a gateway into IM solutions to high-dimensional problems.
@@ 6.2 Nonparametric
At least in principle, one can approach this problem in a manner similar to that in the parametric case above, i.e., via profiling out those nuisance parts of $\mathsf{P}$. Recall that the goal of profiling is to reduce the dimension, so that the compatibility of data with candidate values of the quantity of interest could be directly assessed. Since it is typically the case that the quantity $\Theta$ of interest has some real-world interpretation, there is an opportunity to leverage that interpretation for the aforementioned compatibility assessment without the need of profiling. This is the approach taken in Cella and Martin, (2022), building on the classical work in M-estimation (e.g., Huber, 1981; Stefanski and Boos, 2002) and less-classical work on Gibbs posteriors (e.g., Martin and Syring, 2022; Grünwald and Mehta, 2020; Bissiri et al., 2016; Zhang, 2006), which we summarize briefly below.
@@
The outer supremum, analogous to that in (13), maximizes over all those probability distributions $\mathsf{P}$ such that the relevant feature $\Theta$ of $\mathsf{P}$ takes value $\theta$. This supremum appears because the distribution of $R^{\text{\sc er}}(X^{n},\theta)$ obviously depends on the underlying $\mathsf{P}$, but this is unknown. This puts direct evaluation of the IM contour based on naive Monte Carlo out of reach. Fortunately, validity only requires that certain calibration holds for the one true $\mathsf{P}$, which provides a shortcut. Cella and Martin, (2022) proposed to replace “iid sampling from $\mathsf{P}$ for all $\theta$-compatible $\mathsf{P}$” with iid sampling from the empirical distribution, which is a good estimator of the one true $\mathsf{P}$. This amounts to using the bootstrap (e.g., Efron, 1979; Efron and Tibshirani, 1993; Davison and Hinkley, 1997) to approximate the above contour, and Cella and Martin prove that the corresponding IM is asymptotically valid. Here, we will demonstrate that the proposed variational-like IMs can provide a good approximation for this bootstrap-based contour.
@@
where $\hat{p}$ denotes a kernel density estimate of $p$ based on the observed data. The same settings as in Section 4 were applied, with $\xi$ estimated using $M=200$ Monte Carlo samples, step sizes $w_{t}=2(1+t)^{-1}$, $\alpha=0.1$ and convergence threshold $\varepsilon=0.005.$ Note how the variational approximation is perfect except for small levels of $\alpha$ on the left side, where it is a bit conservative.
@@ 6.3 Semiparametric
Similar to the strategy described above (from Cella and Martin, 2022) for the nonparametric problem, Cahoon and Martin, (2021) proposed a novel strategy wherein a variation on the Kaplan–Meier estimator (e.g., Kaplan and Meier, 1958; Klein and Moeschberger, 2003) is used to obtain a $\widehat{G}$, and then the contour above is approximated by
@@ 7 Conclusion
A number of important and practically relevant extensions to the proposed method can and will be explored. Here we will mention two such extensions. First, recent efforts (e.g., Martin, 2022b ) have focused on incorporating incomplete or partial prior information into the possibilistic IM; something along these lines is sure to be necessary as we scale up to problems of higher dimension. The only downside to the incorporation of partial prior information is that the evaluation of the IM contour is often even more complicated than in the vacuous-prior case considered here. Thanks to the added complexity, efficient numerical approximations are even more important in these partial-prior formulations and, fortunately, we expect that the proposal here will carry over more-or-less directly. Second, there is an obvious question of whether the $\alpha$-specific approximations developed here can, somehow, be stitched together in order to obtain an approximation that is (mostly) insensitive to the specification of any particular $\alpha$ value. Inspired by the recent developments in Jiang et al., (2023), we believe that the answer to this question is Yes, and these important details will be reported elsewhere.

### v3

@@ 4 Beyond basic variational-like IMs
4.1 Setup and proposal
Various tweaks to the basic procedure described in Section 3 might be considered in order applications to speed up computations. If there is a computational bottleneck remaining in the proposal from Section 3, then it would have to be that the IM’s possibility contour must be evaluated at $M$-many points in each iteration of the stochastic approximation. While this is not impractically expensive in some applications, including those presented above, it could be a substantial burden in other applications. A related challenge is that a scalar index $\xi$ on the variational family, as we have focused on exclusively so far, may not be sufficiently flexible. Increasing the flexibility by considering higher-dimensional $\xi$ would likewise increase the computational burden, so care is needed. Here we aim to address both of the aforementioned challenges.
@@ 4.1 Setup and proposal
That is, the components of $\xi$ act as multipliers on the singular values of $J_{n}^{-1}$. This indeed generalizes our previously-considered Gaussian approximation family since, if $\xi$ is a scalar, then we recover the previous covariance matrix: $J_{n}(\xi)=J_{n}/\xi^{2}$. Let $C_{\alpha}^{\xi}(x^{n})$ denote the $100(1-\alpha)$credible set corresponding to the distribution $\mathsf{Q}_{n,\alpha}^{\xi}\in\mathscr{Q}^{\text{var}}$. In the present $d$-dimensional Gaussian case, this is an ellipsoid in $\mathbb{R}^{d}$, i.e.,
$C_{\alpha}^{\xi}(x^{n})=\{\theta\in\mathbb{R}^{d}:(\theta-\hat{\theta}_{n})^{\top}J_{n}(\xi)\,(\theta-\hat{\theta}_{n})\leq\chi_{d}^{2}(1-\alpha)\},$
@@
$C_{\alpha}^{\xi}(x^{n})\supseteq C_{\alpha}(x^{n}),$
@@
$\sup_{\theta\not\in C_{\alpha}^{\xi}(x^{n})}\pi_{n}(\theta)\leq\alpha.$
For the Gaussian family of approximations we are considering here, the mode of $\pi_{n}$, the maximum likelihood estimator, is in the interior of each credible set, so the behavior of $\pi_{n}$ outside $C_{\alpha}^{\xi}(x^{n})$ is determined by its behavior on the boundary. Moreover, since equality in the above display implies a near-perfect match between the IM’s confidence sets and the posited Gaussian credible sets, a reasonable goal is to find a root to the function
$g_{\alpha}(\xi):=\max_{\theta\in\partial C_{\alpha}^{\xi}(x^{n})}\pi_{n}(\theta)-\alpha,$
where $\partial A$ denotes the boundary of a set $A$. In designing an iterative algorithm that makes steps towards this root, it is desirable to allow for steps of different magnitudes depending on the direction, and this requires care. For our present Gaussian approximation family, a natural—albeit minimal—representation of the boundary of the credible region $C_{\alpha}^{\xi}(x^{n})$ is given by the collection of vectors
@@
At least intuitively, $\hat{g}_{\alpha,s}(\xi)$ being less than or greater than 0 determines if the credible set $C_{\alpha}^{\xi}(x^{n})$ is too large or too small, respectively in the $u_{s}$-direction. From here, we can apply the same stochastic approximation procedure to determine a sequence $(\xi^{(t)}:t\geq 1)$, this time of $d$-vectors, i.e.,
@@
that converges to the root of $\hat{g}_{\alpha}$ which we expect to be close to the root of $g_{\alpha}$. The details of this new but familiar strategy are presented in Algorithm 2. The chief advantage of this new strategy compared to that summarized in Algorithm 1 is computational savings: each iteration of stochastic approximation in Algorithm 2 only requires (Monte Carlo) evaluation of the IM contour at $2d$ many points, which would typically be far fewer than then number of Monte Carlo samples $K$ in Algorithm 1 needed to get a good approximation of $\mathsf{Q}_{n,\alpha}^{\xi}\{\pi_{n}(\Theta)>\alpha\}$. That is, the computational complexity of this new strategy is $O(Md)$, where $M$ is the Monte Carlo sample size in (7) to evaluate $\pi_{x^{n}}$.
@@
4.2 Numerical illustrations
We will illustrate this new version of the Gaussian variational family and approximation algorithm with three examples. The first revisits the previous bivariate normal correlation example but with more specific details about timing and accuracy of the approximation; the second is the classical two-parameter gamma model; and the third is a relatively high-dimensional model involving a penalty, which is intended to serve as a gateway into IM solutions to high-dimensional problems.
Example 6.
[…]
  • both strategies have computational complexity linear in $n$, but the naive strategy’s rate of growth is faster than the proposed approximation’s, so the computation time ratio should be roughly constant and greater than 1, and
  • the Gaussian approximation employed in $\hat{\pi}^{\text{approx}}$ improves in accuracy, by the results in Martin and Williams, (2025), so we expect the two contours to merge.
The results in Table 1 confirm these expectations, i.e., the relative computation time holds steady near value 2, so that the naive strategy takes roughly twice as long to compute as the proposed strategy, and the $L_{1}$ distance between the two contours is decreasing with $n$. The computation time comparison is favorable for the naive strategy in this case, since it only involves a scalar parameter, but still the performance difference is substantial.
Sample size, $n$ Relative time $L_{1}$ distance
50 1.92 0.037
100 1.87 0.021
200 1.84 0.011
Table 1: Computation time and accuracy comparisons of the two strategies for evaluating the IM contour in the bivariate normal model: the naive strategy (7) and the proposed approximation as described in Algorithm 2. Based on averages over 100 data sets.
@@ 5.2 Nonparametric
At least in principle, one can approach this problem in a manner similar to that in the parametric case above, i.e., via profiling out those nuisance parts of $\mathsf{P}$. Recall that the goal of profiling is to reduce the dimension, so that the compatibility of data with candidate values of the quantity of interest could be directly assessed. Since it is typically the case that the quantity $\Theta$ of interest has some real-world interpretation, there is an opportunity to leverage that interpretation for the aforementioned compatibility assessment without the need of profiling. This is the approach taken in Cella and Martin, 2022a , building on the classical work in M-estimation (e.g., Huber, 1981; Stefanski and Boos, 2002) and less-classical work on Gibbs posteriors (e.g., Martin and Syring, 2022; Grünwald and Mehta, 2020; Bissiri et al., 2016; Zhang, 2006), which we summarize briefly below.
@@
The outer supremum, analogous to that in (15), maximizes over all those probability distributions $\mathsf{P}$ such that the relevant feature $\Theta$ of $\mathsf{P}$ takes value $\theta$. This supremum appears because the distribution of $R^{\text{\sc er}}(X^{n},\theta)$ obviously depends on the underlying $\mathsf{P}$, but this is unknown. This puts direct evaluation of the IM contour based on naive Monte Carlo out of reach. Fortunately, validity only requires that certain calibration holds for the one true $\mathsf{P}$, which provides a shortcut. Cella and Martin, 2022a proposed to replace “iid sampling from $\mathsf{P}$ for all $\theta$-compatible $\mathsf{P}$” with iid sampling from the empirical distribution, which is a good estimator of the one true $\mathsf{P}$. This amounts to using the bootstrap (e.g., Efron, 1979; Efron and Tibshirani, 1993; Davison and Hinkley, 1997) to approximate the above contour, and Cella and Martin prove that the corresponding IM is asymptotically valid. Here, we will demonstrate that the proposed variational-like IMs can provide a good approximation for this bootstrap-based contour.
@@
where $\hat{p}$ denotes a kernel density estimate of $p$ based on the observed data. The same settings as in Section 3.3 were applied, with $\xi$ estimated using $M=200$ Monte Carlo samples, step sizes $w_{t}=2(1+t)^{-1}$, $\alpha=0.1$ and convergence threshold $\varepsilon=0.005.$ Note how the variational approximation is perfect except for small levels of $\alpha$ on the left side, where it is a bit conservative.
@@ 5.3 Semiparametric
Similar to the strategy described above (from Cella and Martin, 2022a ) for the nonparametric problem, Cahoon and Martin, (2021) proposed a novel strategy wherein a variation on the Kaplan–Meier estimator (e.g., Kaplan and Meier, 1958; Klein and Moeschberger, 2003) is used to obtain a $\widehat{G}$, and then the contour is approximated by
@@ 6 Conclusion
There are a number of limitations associated with what has been proposed here in this paper. Naturally, these limitations lead to open questions and future research directions to be pursued. First, the more sophisticated and efficient of the approximations proposed here, the one in Section 4, is designed specifically for Gaussian variational families. This is not a serious practical limitation since the Gaussian distribution is sure to be a good approximation when $n$ is moderate to large (Martin and Williams, 2025). However, there are surely other variational families for which the density level sets have a clean, more-or-less closed-form expression in terms of the model parameters. Identifying other suitable models that can be used in this efficient approximation will give users more flexibility and ultimately lead to better and more accurate approximations. Second, the proposed approximations depend on specifying a choice of $\alpha$ at the outset, which means that we are effectively only approximating certain features of the possibilistic IM’s contour. Arguably, the holistic nature of IM-based inference suggests that a broad rather than very specific approximation is preferred, and finding a way to stitch together these $\alpha$-specific IM approximations is an important open problem. Inspired by recent developments in Jiang et al., (2024), we believe that the answer to this question is Yes, and these details will be reported elsewhere. Third, the details presented here are specifically focused on the case where prior information about the unknown $\Theta$ is vacuous. Recent efforts (e.g., Martin, 2022b ) have focused on incorporating incomplete or partial prior information into the possibilistic IM construction; something along these lines is sure to be necessary as we scale up to problems of higher dimension. The only downside to the incorporation of partial prior information is that the evaluation of the IM contour is often even more complicated than in the vacuous-prior case considered here. This additional complexity means that efficient numerical approximations are even more important in these partial-prior formulations and, fortunately, we expect that the proposal here will carry over more-or-less directly. Finally, with the exception of Example 8, our focus here was on problems involving relatively low-dimensional unknowns. As is always the case, there are challenges associated with scaling up the proposed approximation strategy. At the moment, we do not know how these challenges will be overcome, but we are inspired by the latest progress presented here and confident that they can be overcome.